OpenStudy (anonymous):

An electric iron has a rating 750 W,220V. Calculate i) the current passing through it ii) its resistance when in use

OpenStudy (radar):

Use the formula W=EI substituting the values provided, 750 = 220 I or I =750/220 = 3.4090909090909090 or approx 3.41 Amps. For II use Ohms Law R=E/I, R = 220/3.41 = 64.53 Ohms. Note that at turn-on, the current will surge as the resistance is low and then increases as the heating element (resistor) heats up.